Express the roots of the equation z3 − α3 = 0 in terms of α and w, where w is
a complex cube root of unity. Use your answer to find the roots of the following
equations in the form a + ib.
1
Expert's answer
2020-05-20T19:10:47-0400

Given,

we have to find the roots of the equation, "z^3 -\alpha^3=0 \ in \ terms\ of \alpha \ and \ \omega ."

We know,

"z^3-\alpha^3 = (z-\alpha)(z^2 + \alpha^2 +z\alpha)=0
\ which \ implies\ z=\alpha\ (a\ root) \ and\ z^2+\alpha^2+z\alpha=0"

Now,

"z^2+\alpha^2+z\alpha=z^2 +2.z.\frac{\alpha}{2}+(\frac{\alpha}{2})^2+\alpha^2-(\frac{\alpha}{2})^2=0
\ implies\ (z+(\frac{\alpha}{2}))^2+(\frac{3\alpha}{4})^2=0"

"(z+(\frac{\alpha}{2}))=\frac{+}{-} \frac{i\sqrt{3}\alpha}{2}"

or, "z=-(\frac{\alpha}{2})\frac{+}{-}\frac{i\sqrt{3}\alpha}{2}"

or, "z=\alpha((\frac{-1}{2})\frac{+}{-}\frac{i\sqrt{3}}{2})"

or, "z=\alpha\omega\ and\ z=\alpha\omega^2\ where \ \omega=((\frac{-1}{2})+\frac{i\sqrt{3}}{2}) \ and \ \omega^2=((\frac{-1}{2})-\frac{i\sqrt{3}}{2})"

Thus, the three roots of the above equation are:

in terms of "\omega" :

"z=\alpha,\alpha\omega\ and \ \alpha\omega^2"

in terms of a+ib:

"z=\alpha,\alpha" "((\frac{-1}{2})+\frac{i\sqrt{3}}{2}) \ and\ \alpha((\frac{-1}{2})-\frac{i\sqrt{3}}{2})".